OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9

OpenStudy (australopithecus):

you would substitute the value you got for x into one of the original equations

OpenStudy (australopithecus):

to solve for y

OpenStudy (australopithecus):

by solving for y and substituting it into the other equation you are elminating the y variable from that equation and therefore left with an equation with only x as the variable

OpenStudy (australopithecus):

once you have a value of x you can just sub that into either of the equations to solve for y

OpenStudy (australopithecus):

system of equations is just a way to see where two functions intersect

OpenStudy (australopithecus):

so if you had two equations that when graphed formed a line and you used system of equations with them you could find out at what point they meet or if they ever meet

OpenStudy (australopithecus):

and by meet i mean share a common point

OpenStudy (australopithecus):

hope that helps

OpenStudy (anonymous):

i got x =2+y

OpenStudy (australopithecus):

I confused the question sorry

OpenStudy (australopithecus):

4x − y = 8 4x - y - 8 + y = 8 - 8 + y 4x - 8 = y so we sub into the equation 4x - 8 to replace the y variable so, 2x + 3y = 9 becomes 2x + 3(4x-8) = 9

OpenStudy (anonymous):

thank you i got it now

OpenStudy (australopithecus):

do you understand the theory behind this now as well?

OpenStudy (anonymous):

yes i do

OpenStudy (australopithecus):

here is a picture of where they intersect, note there are 3 solutions to these type of problems -where you get 1 or more points so you will just get possible values of x and y as a solution -where you get no solution variables cancel out and you are left with an inequailty that doesnt exist like 2 = 4 or something -Lastly if you get something like 8x = 8x (the 8 cancels out) or x = x, or y=y both equations are identical and therefore overlap (they are the same graph so they share the same points always)

OpenStudy (anonymous):

ii understand

OpenStudy (australopithecus):

an example of the second options would be something like graph y = 8x - 1 and graph y = 8x + 1

OpenStudy (australopithecus):

option that shouldn't be pluralized

OpenStudy (australopithecus):

if you try to solve that system of equation for y = 8x +1 and y=8x-1 you will see that they dont intersect you will get something weird

OpenStudy (australopithecus):

but yeah hope this helps i will stop blathering on now :)

OpenStudy (australopithecus):

and by something weird I mean +1 = -1 which doesn't make sense

OpenStudy (anonymous):

ty